NEW BANDS OF NICKEL FLUORIDE IN THE NEAR INFRARED BY INTRACAVITY LASER ABSORP-TION SPECTROSCOPY

\underline{LEAH C. O’BRIEN}, KIMBERLY HANDLER, Department of Chemistry, Southern Illinois University, Edwardsville, IL 62026-1652; JAMES J. O’BRIEN, Department of Chemistry and Biochemistry, University of Missouri, St Louis, MO 63121-4499.

Several new electronic transitions of NiF with red-degraded bandheads near 12896 and 13497 cm^{−1}were recorded and analyzed. All bands have a common lower level: namely, v=0 of the A ^{2}Δ_{5/2} state. The observed ^{58}NiF and ^{60}NiF isotopologue splitting in each band suggests tentative assignments as the (3,0) and (2,0) vibrational bands of the [12.6]Ω=5/2 - A ^{2}Δ_{5/2} transition and the (2,0) band of a new [13.1]Ω=5/2 - A ^{2}Δ_{5/2} transition of NiF. Results of the analysis will be presented. The gas phase NiF molecules were produced using a nickel-lined hollow cathode in an argon-based electric discharge with a small amount of SF_{6}.